# LaMoD: Latent Motion Diffusion Model For Myocardial Strain Generation

### v1

@@ Abstract
Motion and deformation analysis of cardiac magnetic resonance (CMR) imaging videos is crucial for assessing myocardial strain of patients with abnormal heart functions. Recent advances in deep learning-based image registration algorithms have shown promising results in predicting motion fields from routinely acquired CMR sequences. However, their accuracy often diminishes in regions with subtle appearance change, with errors propagating over time. Advanced imaging techniques, such as displacement encoding with stimulated echoes (DENSE) CMR, offer highly accurate and reproducible motion data but require additional image acquisition, which poses challenges in busy clinical flows. In this paper, we introduce a novel Latent Motion Diffusion model (LaMoD) to predict highly accurate DENSE motions from standard CMR videos. More specifically, our method first employs an encoder from a pre-trained registration network that learns latent motion features (also considered as deformation-based shape features) from image sequences. Supervised by the ground-truth motion provided by DENSE, LaMoD then leverages a probabilistic latent diffusion model to reconstruct accurate motion from these extracted features. Experimental results demonstrate that our proposed method, LaMoD, significantly improves the accuracy of motion analysis in standard CMR images; hence improving myocardial strain analysis in clinical settings for cardiac patients. Our code will be publicly available on upon acceptance.
@@ 1 Introduction
Motion and deformation analysis of CMR videos provides valuable measurements to quantify myocardial strain in patients with heart diseases, offering clinically significant data for disease assessment, diagnosis, and treatment planning [16, 3, 23, 28, 29]. CMR feature tracking (FT) is widely used in clinical practice to evaluate myocardial motion, deformation, and strain functions [24, 14, 17, 18]. Such a technique employs optical flow-based algorithms [10] to track image features or patterns within the myocardium throughout the cardiac cycle. While FT is convenient and integrates well into routine clinical workflows, this technique is generally less accurate due to its limited motion tracking accuracy [1, 31].
With recent advancements in deep learning, several research groups have utilized image registration-based networks to predict myocardial motion and strain from CMR images [14, 17, 30]. Existing methods incorporated new regularization techniques, such as enforcing temporal consistency [17], or introducing the prior knowledge of cardiac biomechanics  [20, 34, 19], to achieve improved motion prediction quality. However, these approaches still partially address the challenges of detecting motion in regions with subtle image appearance changes (i.e., the myocardium mid-wall motion) [1, 6, 15], leading to compromised accuracy of myocardial strain measurements. To alleviate this issue, another research line has utilized “ground-truth” data from the advanced imaging technique DENSE, which provides highly accurate and reproducible myocardial motion data to supervise the motion learning process [26, 27]. In contrast to conventional CMR techniques, DENSE directly encodes tissue displacement within the imaging data, allowing for precise quantification of myocardial motion throughout the cardiac cycle. Despite promising progress in predicting myocardial motion for improved strain analysis under the guidance of DENSE [26, 27], current approaches heavily rely on features extracted from segmented myocardial contours rather than the underlying cardiac motions, potentially compromising strain accuracy. Furthermore, the network training involved only DENSE contours that may not fully generalize to standard CMR images.
@@
    To the best of our knowledge, our method is the first to leverage latent diffusion models in the motion feature space to produce precise myocardial strain that is comparable to DENSE from standard CMR videos. This approach opens promising research avenues for transferring knowledge from advanced strain imaging to routinely acquired CMR data, maximizing benefits for patients with cardiac diseases.
@@
    Our experimental results demonstrate a substantial improvement in motion detection and strain generation accuracy compared to existing approaches. Future work will involve meticulous validation of model generalizability as additional patient data becomes available.
@@ 3.2 Diffusion Model In Latent Motion Spaces
Inspired by the Denoising Diffusion Probabilistic Models (DDPM)[9, 21], we formulate the model as a latent Markov chain with $M$ steps. The framework has a forward and reverse diffusion process. The forward process slowly distorts the input latent features over $m\in[1,2,...,M]$ steps using the smoothed version of a randomly sampled Gaussian noise, i.e. $\epsilon^{\prime}=\mathcal{K}(\epsilon)$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, $\mathcal{K}(.)$ is a Gaussian smoothing kernel and $\mathcal{N}$ defines a normal distribution.
@@
1:  $\{\boldsymbol{z}^{(0)}\}=\{z^{1},\dots,z^{T}\}=\mathcal{E}_{R}(\{I^{\tau}\})$ Diffusion Model Inference
2:  $m^{\prime}\sim\text{Uniform}(\{1,\dots,M\})$
3:  $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
4:  $\boldsymbol{z}^{(m^{\prime})}=\sqrt{\bar{\alpha_{m^{\prime}}}}\boldsymbol{z}^{(0)}+\sqrt{1-\bar{\alpha_{m^{\prime}}}}~{}\mathcal{K}(\epsilon)$
5:  for $m=m^{\prime},m^{\prime}-1,\dots,1$ do
6:     $\boldsymbol{\hat{z}}^{(m-1)}=\frac{1}{\sqrt{\alpha_{m}}}(\boldsymbol{\hat{z}}^{(m)}-\frac{1-\alpha_{m}}{\sqrt{1-\bar{\alpha_{m}}}}\epsilon_{\theta}(\boldsymbol{\hat{z}}^{(m)},m))+\sigma_{m}\mathcal{K}(\gamma)$
7:  end for
8:  $\{\hat{\Phi}^{\tau}\}=\mathcal{D}_{\eta}(\boldsymbol{\hat{z}}^{(0)})$;
9:  return  $\{\hat{\Phi}^{\tau}\}$
@@ 4 Experimental Evaluation
The training of all our experiments is implemented on an server with AMD EPYC $7502$ CPU of $126$GB memory and Nvidia GTX $3090$Ti GPUs. We train our networks using the Adam optimizer [11] with maximal $2000$ epochs with early stop strategy. The batch size is set to $32$ and weight decay weights are set to $\lambda_{\epsilon}=\lambda_{M}=1E-4$. The hyper-parameters are optimized with grid search strategy. The optimal learning rate is $41E-$ 4 and the optimal loss weight $\alpha=1E-2$.
@@ 4.1 Dataset
Standard cine CMR videos. We carefully selected $105$ short-axis cine CMRs slices of $40$ subjects from the DENSE dataset mentioned above, including $14$ patients and $26$ volunteers. All scans were acquired during repeated breath holds covering the left ventricle (LV) (field of view, 320 × 320 to 380 × 380 mm2; temporal resolution, 30–55 msec, depending on heart rate). Each selected cine slice corresponds to a DENSE scan from the same patient at the same spatial location (within ±$\pm 2\text{mm}$), allowing the DENSE displacement field to serve as the ground-truth motion for the cine slices.
@@ 4.2 Experimental Design
We used all standard CMR and DENSE videos to train our registration network that can effectively learn latent motion features. However, we only use DENSE motion to train the diffusion model to reconstruct the myocardial deformations. In particular, we split the DENSE data as $538$ for training, $101$ for validation, and $102$ for testing. The paired DENSE data with cine CMR videos were used for regional segmental strain comparison in our experimental evaluation.
@@
We then generate strain maps from the predicted motion fields for DENSE input and compare it with all baseline algorithms.We performed a quantitative analysis of strain errors on regional segmental strains. Consistent with previous studies [26, 27], we divided the myocardium in each DENSE slice into six segments, starting from the right ventricle insertion point and proceeding counterclockwise. We then calculated the average absolute error for strain evaluation in each segment.
Evaluate strain error on standard cine CMR videos. For cine sequences, our evaluation focuses exclusively on segmental strain error due to the significant differences between the paired DENSE contours and the collected cine CMR myocardium contours. This mainly due to the spatial misalignment between the myocardium regions from input cine images and the DENSE-derived ground truth, which are collected from separate scans, making pixel-wise error computation impractical. We apply the same segmental strain approach as used for DENSE data, dividing the myocardium into six segments and calculating the average absolute strain error for each. Additionally, we compare the strain values with those obtained using widely used commercial software (SuiteHeart version 5.0.4; NeoSoft) based on Feature Tracking (FT) in clinical settings.
@@ 4.3 Experimental Results
Fig. 2 (test on DENSE data) presents the visualizations of the predicted end-systolic displacement fields and the circumferential (Ecc) maps for all models compared to the ground truth from DENSE ground truth. To provide a comprehensive evaluation, we include examples from both healthy volunteers (top panel: ground A) and patients with heart failure (bottom panel: ground B), specifically those with left bundle branch block (LBBB). The predictions generated by our method consistently show a closer resemblance to the ground truth across all cases. These results collectively demonstrate that our method provides more accurate and robust performance compared to the baseline models.
Figure 2: Left to right: examples of end-systolic displacement and circumferential (Ecc) strain maps (contraction in blue vs. stretch in red) from DENSE ground truth and predictions from our model vs. other baseline algorithms. Top to bottom: motion and strain from (A) a healthy volunteer and (B) heart failure patient with left bundle branch block.
Figure 3: Left to right: examples of end-systolic displacement and circumferential (Ecc) strain maps from standard cine CMRs and predictions from our model vs. other baseline algorithms. Top to bottom: motion and strain from (A) a healthy volunteer and (B) heart failure patient with left bundle branch block.
@@
Similarly, the Fig. 3 (test on standard cine CMRs) visualizes the predicted motion fields and the circumferential (Ecc) maps for all models compared to paired DENSE dataset. Note that the DENSE myocardium contours are slightly different from cine CMRs due to a different scanning time.
Fig. 4 displays a quantitative comparison of displacement field EPE error and segmental circumferential strain error between our model and all baselines. The left two panels of Fig. 4 demonstrate the testing results on DENSE dataset, indicating that our method significantly outperforms the baselines in terms of both displacement error and circumferential strain error. The right panel of Fig. 4 shows the testing results on the cine CMRs dataset. The consistently superior performance of our method suggests that our model, LaMoD, is able to achive improved quality of strain values from standard cine CMR videos when compared to all baseline methods.
@@ 5 Conclusion
In this paper, we introduced LaMoD, a novel Latent Motion Diffusion model that significantly advances the field of cardiac motion analysis by predicting accurate motions/strain from standard CMR videos. Our approach effectively addresses the challenges of detecting subtle myocardial movements, particularly in intramyocardial regions, by leveraging a pre-trained registration network and a probabilistic latent diffusion model guided by DENSE CMRs. Experimental results demonstrate that LaMoD outperforms existing methods in motion prediction and strain generation accuracy, bridging the gap between routinely acquired CMR sequences and advanced DENSE imaging. This advancement has the potential to improve cardiac disease assessment and treatment planning in clinical settings without requiring additional DENSE scans, ultimately improving patient care. Future work will focus on validating the model’s generalizability across diverse patient populations and cardiac conditions.

### v2

@@ Abstract
Motion and deformation analysis of cardiac magnetic resonance (CMR) imaging videos is crucial for assessing myocardial strain of patients with abnormal heart functions. Recent advances in deep learning-based image registration algorithms have shown promising results in predicting motion fields from routinely acquired CMR sequences. However, their accuracy often diminishes in regions with subtle appearance changes, with errors propagating over time. Advanced imaging techniques, such as displacement encoding with stimulated echoes (DENSE) CMR, offer highly accurate and reproducible motion data but require additional image acquisition, which poses challenges in busy clinical flows. In this paper, we introduce a novel Latent Motion Diffusion model (LaMoD) to predict highly accurate DENSE motions from standard CMR videos. More specifically, our method first employs an encoder from a pre-trained registration network that learns latent motion features (also considered as deformation-based shape features) from image sequences. Supervised by the ground-truth motion provided by DENSE, LaMoD then leverages a probabilistic latent diffusion model to reconstruct accurate motion from these extracted features. Experimental results demonstrate that our proposed method, LaMoD, significantly improves the accuracy of motion analysis in standard CMR images; hence improving myocardial strain analysis in clinical settings for cardiac patients. Our code is publicly available at https://github.com/jr-xing/LaMoD.
@@ 1 Introduction
Motion and deformation analysis of CMR videos provides valuable measurements to quantify myocardial strain in patients with heart disease, offering clinically significant data for disease assessment, diagnosis, and treatment planning [17, 3, 24, 29, 30]. CMR feature tracking (FT) is widely used in clinical practice to evaluate myocardial motion, deformation, and strain functions [25, 15, 18, 19]. Such a technique employs optical flow-based algorithms [10] to track image features or patterns within the myocardium throughout the cardiac cycle. While FT is convenient and integrates well into routine clinical workflows, this technique is generally less accurate due to its limited motion tracking accuracy [1, 32].
With recent advancements in deep learning, several research groups have utilized image registration-based networks to predict myocardial motion and strain from CMR images [15, 18, 31]. Existing methods incorporated new regularization techniques, such as enforcing temporal consistency [18], or introducing prior knowledge of cardiac biomechanics  [21, 35, 20], to achieve improved motion prediction quality. However, these approaches still partially address the challenges of detecting motion in regions with subtle image appearance changes (i.e., the myocardium mid-wall motion) [1, 6, 16], leading to compromised accuracy of myocardial strain measurements. To alleviate this issue, another research line has utilized “ground-truth” data from the advanced imaging technique DENSE, which provides highly accurate and reproducible myocardial motion data to supervise the motion learning process [27, 28]. In contrast to conventional CMR techniques, DENSE directly encodes tissue displacement within the imaging data, allowing for precise quantification of myocardial motion throughout the cardiac cycle. Despite promising progress in predicting myocardial motion for improved strain analysis under the guidance of DENSE [27, 28], current approaches heavily rely on features extracted from segmented myocardial contours rather than the underlying cardiac motions, potentially compromising strain accuracy. Furthermore, the network training involved only DENSE contours that may not fully generalize to standard CMR images.
@@
    Our method is the first to leverage latent diffusion models in the motion space to produce highly accurate myocardial strain from standard CMR videos.
@@
    Opens promising research avenues for transferring knowledge from advanced strain imaging to routinely acquired CMR data; hence maximizing benefits for patients with cardiac diseases.
@@ 3.2 Diffusion Model In Latent Motion Spaces
Inspired by the Denoising Diffusion Probabilistic Models (DDPM)[9, 22], we formulate the model as a latent Markov chain with $M$ steps. The framework has a forward and reverse diffusion process. The forward process iteratively adds random Gaussian noise to the input latent features over $m\in[1,2,...,M]$ steps. Similar to [11], we employ smoothed Gaussian noise $\epsilon^{\prime}$ in the diffusion process to facilitate faster optimization convergence compared to normal Gaussian noise $\epsilon$, i.e. $\epsilon^{\prime}=\mathcal{K}(\epsilon)$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ and $\mathcal{K}(\cdot)$ is a Gaussian smoothing kernel.
@@
1:  $\boldsymbol{z}^{(0)}=\{z^{1},\dots,z^{T}\}=\mathcal{E}_{R}(\{I^{0},\dots,I^{T}\})$ Diffusion Model Inference
2:  $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
3:  $\boldsymbol{z}^{(M)}=\sqrt{\bar{\alpha_{M}}}\boldsymbol{z}^{(0)}+\sqrt{1-\bar{\alpha_{M}}}~{}\mathcal{K}(\epsilon)$
4:  for $m=M,M-1,\dots,1$ do
5:     $\boldsymbol{\hat{z}}^{(m-1)}=\frac{1}{\sqrt{\alpha_{m}}}(\boldsymbol{\hat{z}}^{(m)}-\frac{1-\alpha_{m}}{\sqrt{1-\bar{\alpha_{m}}}}\epsilon_{\theta}(\boldsymbol{\hat{z}}^{(m)},m))+\sigma_{m}\mathcal{K}(\gamma)$
6:  end for
7:  $\{\hat{\Phi}^{\tau}\}=\mathcal{D}_{\eta}(\boldsymbol{\hat{z}}^{(0)})$;
8:  return  $\{\hat{\Phi}^{\tau}\}$
@@ 4 Experimental Evaluation
The training of all our experiments is implemented on an server with AMD EPYC $7502$ CPU of $126$GB memory and Nvidia GTX $3090$Ti GPUs. We train our networks using Adam optimizer [12] with maximal $2000$ epochs with the early stop strategy. The batch size is set to $32$ and weight decay weights are set to $\lambda_{\epsilon}=\lambda_{M}=1E-4$. The hyper-parameters are optimized with grid search strategy. The optimal learning rate is $41E-$ 4 and the optimal loss weight is $\alpha=1E-2$.
@@ 4.1 Dataset
Standard cine CMR videos. We tested our proposed model on $105$ short-axis cine CMRs slices of $40$ subjects from the DENSE dataset mentioned above, including $14$ patients and $26$ volunteers. All scans were acquired during repeated breath hold covering the left ventricle (LV) (field of view, 320 × 320 to 380 × 380 mm2; temporal resolution, 30–55 msec, depending on heart rate). Each selected cine slice corresponds to a DENSE scan of the same patient at the same spatial location (within ±$\pm 2\text{mm}$), allowing the DENSE displacement field to serve as the ground-truth motion for the cine slices.
@@ 4.2 Experimental Design
We used all standard CMR and DENSE videos to train our registration network that can effectively learn latent motion features. However, we only use DENSE motion to train the diffusion model to reconstruct myocardial deformations. In particular, the DENSE data set was divided into 538 samples for training, 101 for validation and 102 for testing. We employed a site-balanced splitting strategy, maintaining consistent proportions from each site across all sets to ensure representative sampling and mitigate site-specific biases. Paired DENSE data with cine CMR videos were used for regional segmental strain comparison in our experimental evaluation.
@@
We then generate strain maps from the predicted motion fields for the DENSE input and compare it with all baseline algorithms. We performed a quantitative analysis of strain errors on regional segmental strains. Consistent with previous studies [27, 28], we divided the myocardium in each DENSE slice into six segments, starting from the right ventricle insertion point and proceeding counterclockwise. We then calculated the average absolute error for strain evaluation in each segment.
Evaluate strain error on standard cine CMR videos. For cine sequences, our evaluation focuses exclusively on segmental strain error due to the significant differences between the paired DENSE contours and the collected cine CMR myocardium contours. This is mainly due to the spatial misalignment between the myocardium regions from input cine images and the DENSE-derived ground truth, which are collected from separate scans, making pixel-wise error computation impractical. We apply the same segmental strain approach as used for the DENSE data, dividing the myocardium into six segments and calculating the average absolute strain error for each. Furthermore, we compared strain values with those obtained using widely used commercial software (SuiteHeart version 5.0.4; NeoSoft) based on Feature Tracking (FT) in clinical settings.
@@ 4.3 Experimental Results
Figure 2: Panel A (healthy volunteer) & B (heart failure patient with left bundle branch block): exemplary comparison of end-systolic displacement and circumferential strain maps (Ecc) derived from DENSE input across all methods. Left to right: DENSE ground truth and predictions from our model vs. baselines. Top to bottom: enlarged view of selected displacement region; full displacements; circumferential strain maps (contraction in blue vs. stretch in red).
Fig. 2 (test on DENSE data) presents the visualizations of the predicted end-systolic displacement fields and the circumferential (Ecc) strain maps for all models compared to the ground truth from DENSE ground truth. To provide a comprehensive evaluation, we include examples from both healthy volunteers (top panel: ground A) and patients with heart failure (bottom panel: ground B), specifically those with left bundle branch block (LBBB). The predictions generated by our method consistently show a closer resemblance to the ground truth across all cases. These results collectively demonstrate that our method provides more accurate and robust performance compared to the baseline models.
Figure 3: Panel A (healthy volunteer) & B (heart failure patient with left bundle branch block): exemplary comparison of end-systolic displacement and circumferential strain maps (Ecc) derived from standard cine MRI videos.
@@
Similarly, the Fig. 3 (test on standard cine CMRs) visualizes the predicted motion fields and the circumferential (Ecc) strain maps for all models compared to paired DENSE dataset. Note that the DENSE myocardium contours are slightly different from cine CMRs due to a different scanning time.
Fig. 4 displays a quantitative comparison of displacement field EPE error and segmental circumferential strain error between our model and all baselines. The left two panels demonstrate the testing results on DENSE dataset, indicating that our method significantly outperforms the baselines in terms of both displacement and circumferential strain error. The right panel shows the testing results on the cine CMRs dataset. It shows that our method consistently achieves superior performance of myocardial strain quality over all baselines.
Figure 5: A comparison of time-sequential myocardial strain generated by DENSE vs. our model throughout the cardiac cycle.
Fig.5 illustrates a comparison between circumferential strain computed from DENSE and our predictions across evenly sampled time frames throughout the cardiac cycle. The visual similarity between the top and bottom rows demonstrates that our method effectively captures the strain patterns at various phases of cardiac motion.
@@ 5 Conclusion
In this paper, we introduced LaMoD, a novel Latent Motion Diffusion model that predicts highly accurate motions/strain from standard CMR videos. Our approach effectively addresses the challenges of detecting subtle myocardial movements, particularly in intramyocardial regions, by leveraging a pre-trained registration network and a probabilistic latent diffusion model in the latent motion space guided by DENSE CMRs. Experimental results demonstrate that LaMoD outperforms existing methods in motion prediction and strain generation accuracy. Our work has great potential to improve cardiac disease assessment based on strain, as well as treatment planning without requiring additional DENSE scans; hence ultimately improving patient care. Future work will focus on further validating the model’s generalizability across diverse patient populations and clinical environments.
@@
Disclosure of Interests. The authors have no competing interests to declare that are relevant to the content of this article.